# Modelling bulk mechanical effects in a planar cellular monolayer

### v1

@@ III.1 Regular hexagonal cells: parameter variation
$-\Gamma_{A}$ controls the strength of cell-cell adhesion, so that increasing $-\Gamma_{A}$ acts to increase lateral face area. Accordingly, decreasing $-\Gamma_{A}$ penalises contact to the extent that $H$ falls to zero at $-\Gamma_{A}\approx-0.38$ (Figure 4). While the cell height monotonically increases with $-\Gamma_{A}$, there is a maximum in apical area and perimeter at $-\Gamma_{A}\approx-0.25$. Strong adhesion elongates cells but also drives them into a floppy regime. This can be understood from the terms $\tfrac{1}{2}\Gamma_{A}H-2\Gamma_{L}L_{0}$ in $T^{2D}$ in (14c): increasing $-\Gamma_{A}$ has an effect roughly equivalent to increasing $L_{0}$, which as already been seen to induce loss of rigidity. Once in the floppy regime, the asymptote (50) is recovered, with $H\propto-\Gamma_{A}$, $A\propto-\Gamma_{A}^{-1}$ and $L\propto-\Gamma_{A}^{0}$, consistent with very strong adhesion driving the formation of columnar cells. We will show below how columnar cells can also remain rigid for large $-\Gamma_{A}$.
@@
Figure 4: Impact of varying $-\Gamma_{A}$, presented following the format of Fig. 2. (b) The base case cell shape at $\Gamma_{A}=-0.5$ is shown in grey (solid symbol); cells for $\Gamma_{A}={0.25,-1.0}$ are marked by open symbols. In all panels, $\Gamma_{a}=\Gamma_{L}=0.1,L_{0}=2.0$. The dashed lines in (a,c) show the asymptote (50) as $-\Gamma_{A}\rightarrow\infty$.
@@
Figure 5: Impact of varying $\Gamma_{L}$, presented following the format of Fig. 2. (b) The base case cell shape at $\Gamma_{L}=0.1$ is shown in grey (solid symbol); cells for $\Gamma_{L}={0.025,0.25}$ are marked by open symbols. In all panels, $\Gamma_{a}=0.1,\Gamma_{A}=-0.5,L_{0}=2.0$ The dashed lines show the asymptote as $\Gamma_{L}\rightarrow\infty$ as calculated in (47).
@@
The transitions described above are summarized in Fig. 6, which shows six 2D slices through 4D parameter space, for hexagonal cells. In comparison to the standard 2D model, there are multiple paths in parameter space along which the system loses rigidity, or along which rigid solutions become unphysical. The $(L_{0},\Gamma_{L})$ plane (Fig. 6c) is the most direct analogue of the standard 2D model (Fig. 10), but illustrates significant variation in the value of $L_{0}$ at which the rigidity transition takes place. Figure 6(a) offers an example where rigid solutions are lost either because the apical area vanishes (by decreasing $L_{0}$ from the baseline value, for fixed $-\Gamma_{A}$), or because of the termination of a steady solution branch at a saddle-node bifurcation (by simultaneous reduction of $L_{0}$ and $-\Gamma_{A}$), at which the eigenvalue corresponding to the dilational mode of the Hessian vanishes. Figure 6(e) shows how variation of both $-\Gamma_{A}$ and $\Gamma_{a}$ can lead to loss of rigidity (increasing $-\Gamma_{A}$ from the baseline case), flattening of the cell (reduction of $-\Gamma_{A}$) or cell elongation (via simultaneous increase of $-\Gamma_{A}$ and $\Gamma_{a}$). Fig. 6(d,e) demonstrates distinct floppy states, either squamous (low $\Gamma_{a}$) or columnar (high $\Gamma_{a}$); these are reflected in the two branches of the rigid/floppy boundary in Fig. 6(a,b).
@@
We used asymptotic approaches (Appendix E) to identify the limiting location of the rigid/floppy transition, such as the limit $\Gamma_{a}\approx\Gamma_{L}L_{0}$ (51) for large $-\Gamma_{A}$ (Fig. 6a,e,f), (45) for large $\Gamma_{a}$ (Fig. 6b,d,e) and (47) for large $\Gamma_{L}$ (Fig. 6c,d,f). The boundary at which rigid cells become strongly elongated in Fig. 6(b) is $L_{0}\approx-\sqrt{3}/(48\Gamma_{L})$ (46b)3. Similarly, rigid cells become strongly flattened at the threshold (48) for large $\Gamma_{L}$ (Fig. 6(f); the boundary turns out to be almost independent of $\Gamma_{L}$) and columnar at $\Gamma_{a}\approx-L_{0}^{3}/192$ (49) (Fig. 6c). These limiting approximations reflect dominant physical balances. Cells can remain rigid for large $-\Gamma_{A}$ when $\Gamma_{a}$ is also large (Fig. 6e), with the associated tensions balancing to determine $H\approx\sqrt{3}\Gamma_{A}^{2}/(96\Gamma_{a}^{2})$ (52). In none of these limits does the 3D model recover results from the classical 2D model.
@@
Figure 6: Cross-sections of $(L_{0},\Gamma_{a},-\Gamma_{A},\Gamma_{L})$-parameter space, intersecting the base case (black circle). In each plot, only two parameters are varied. Hatching shows regions where there are no physical solutions. Grey regions show the floppy regime. Blue lines show where $T^{2D}=0$, corresponding to a rigidity transition; blue dashed lines show asymptotes at $\Gamma_{a}\rightarrow\infty,\Gamma_{A}\rightarrow-\infty$ and $\Gamma_{L}\rightarrow\infty$, calculated using equations (45), (51) and (47) respectively. Black dashed lines show asymptotes marking where $H\rightarrow 0$ (for large $\Gamma_{L}$, using (48), in (f)) or where $H\rightarrow\infty$ (for large $\Gamma_{a}$, using (46b), in (b, e); and for large $\Gamma_{L}$, using (49, in (c)).
@@ III.2 Regular hexagonal cells: external loading
The slope of the volume curves in Fig. 7(c) illustrate the cell bulk modulus under in-plane and isotropic compression. The full stiffness matrix is evaluated in Appendix H. Second derivatives of the energy in (60) illustrate how the coupling between height, perimeter and area variations in the present 3D formulation contribute to stiffness. The first derivatives of the energy in (60) illustrate the contribution of prestress.
@@ III.3 Disordered Monolayers
Figure 8: Distributions of cell height perturbations $H^{(1)}$ evaluted using (44), for three different monolayers using base-case parameters, $L_{0}=2,\Gamma_{a}=0.1,\Gamma_{A}=-0.5,\Gamma_{L}=0.1$.
@@
Figure 9: Simulations of a monolayer with $N_{c}=500$ cells and $N_{v}=1080$ apical vertices, successively relaxed through values of $L_{0}$. (a) Distribution of the shape parameter ${L_{h}}/{\sqrt{A_{h}}}$ for varying $L_{0}$ with $\Gamma_{a}=0.1,\Gamma_{A}=-0.5,\Gamma_{L}=0.1$, coloured by the number of apical edges of each cell. White circles indicate outliers and the whiskers show the data extent excluding outliers, while the box shows the interquartile range. The red dots indicate the value of the shape parameter for an array of hexagons, which here is above the value at transition of $s_{6}\approx 3.72$. Thresholds $s_{4}=4$ and $s_{5}\approx 3.81$ (see (18)) are shown with dashed lines. (b) The eigenvalues of the Hessian for the monolayers show in (a), for $L_{0}=\{2.85,2.9,2.95,3.0\}$. (c) Number of connected components $N_{cc}^{\epsilon}$ in each monolayer (blue squares, left axis) and number of vertices in the largest connected component $N^{\varepsilon}_{\max}$ (red diamonds, right axis) as $L_{0}$ is varied. (d) Monolayers for $L_{0}=\{2.88,2.89,2.9,2.95\}$ showing edges with tension above a threshold value of $\varepsilon=0.0005$ in red.
@@ IV Discussion
We have shown how a planar epithelium, modelled using a mechanical energy $U$ that reflects the 3D structure of cells, can be represented in a 2D formulation. To achieve this, we introduced a constraint (42) that restricted height variations between neighbouring cells. A global force balance was used to determine the height $H$ of the cell population. This, together with the involvement of an energy that penalises deviations of the total cell area from a target value, leads to a 2D model that is distinct from the classical formulation in which apical area $A$ and perimeter $L$ contribute separately to $U$. The coupling of $A$, $H$ and $L$ leads to a more intricately-structured parameter space (Fig. 6) that reveals multiple pathways to rigidity transitions, regulated by variation of four independent dimensionless constitutive parameters. This suggests that a monolayer changing state, for example in a wound-healing response, need not rely wholly on the cortical actin structures at the cell edge. This is consistent with prior models [10] demonstrating regulation of rigidity via variations of the bulk target area through, for example, the parameter $a_{0}^{*}/(V_{0}^{*})^{2/3}$. ($a_{0}$ was scaled out of the present model using (13); decreasing $a_{0}$ relative to the baseline case causes $L_{0}$ to increase and $-\Gamma_{A}$, $\Gamma_{a}$ and $\Gamma_{L}$ to fall, with $L_{0}(-\Gamma_{A})^{2}$, $\Gamma_{a}L_{0}$, $\Gamma_{L}L_{0}^{2}$, $\Gamma_{A}^{2}/\Gamma_{L}$, $\Gamma_{a}/(-\Gamma_{A})^{2}$ and $\Gamma_{L}/(-\Gamma_{A})^{4}$ held constant; this path in parameter space does not lie in any of the planes shown in Fig. 6.)
@@ Appendix A Incidence and adjacency matrices
Assigning orientations to all quantities, signed incidence matrices $\mathsf{A}$, $\mathsf{B}$ and $\mathsf{C}$ with components $C_{hi}=\{\mathsf{C}\}_{hi}$, $B_{ij}=\{\mathsf{B}\}_{ij}$ and $A_{jk}=\{\mathsf{A}\}_{jk}$ define topological arrangements. Specifically, an edge orientation is defined by a tangent with magnitude equal to the edge length; a face orientation is defined by a normal with magnitude equal to the face area; $C_{hi}=1$ ($-1$) if the normal to face $i$ points into (out of) cell $h$; $B_{ij}=1$ ($-1$) if edge $j$ bounds face $i$ and is congruent (or anticongruent) to its orientation in a right-handed sense; $A_{jk}=1$ ($-1$) if edge $j$ points into (out of) vertex $k$. The components of $\mathsf{A}$, $\mathsf{B}$ and $\mathsf{C}$ are zero otherwise.
@@
The factor $\tfrac{1}{6}$ in $\mathsf{D}^{\mathrm{cv}}$ accounts for the fact that each vertex $k$ of cell $h$ neighbours three edges of the cell and each edge bounds two faces of the cell. For the prismatic cells illustrated Fig. 1, distinguishing apical ($+$), lateral $(l)$ and basal ()(-)( - ) objects, the adjacency matrices inherit the structure of the incidence matrices via
@@ Appendix B Geometric identities
$L_{h}^{+}$ $=\textstyle{\sum_{i,j}}|C_{hi}^{l}||B_{ij}^{l+}||\mathbf{t}^{+}_{j}|,$ (38a)
$A_{h}^{+}$ $=\tfrac{1}{2}{\textstyle{\sum_{i,j}}}|C_{hi}^{+}|B_{ij}^{++}\mathbf{c}_{j}^{+}\cdot(\mathbf{t}_{j}^{+}\times\hat{\mathbf{n}}^{+}_{hi}),$ (38b)
@@
Rigid hexagons satisfy (16a,b); floppy solutions satisfy $A_{h}=1$, $L_{h}=L_{0}^{2D}$, as illustrated in Fig. 10(b,c) (replicating [1], for example), which can be compared to figures 2. The parameter space map in Fig. 10(a) can be compared to Fig. 6(c). The rigid-floppy transition lies at $L_{0}^{2D}=s_{6}\approx 3.72$. Rigid solutions collapse to zero area along $L_{0}^{2D}=0$ for $\Gamma^{2D}>\Gamma^{*}\equiv\sqrt{3}/12$. Rigid solution branches terminate via a saddle-node bifurcation at $L_{0}^{2D}=-({8\sqrt{2}}/(3\Gamma))\left(\Gamma^{*}-\Gamma^{2D}\right)^{\tfrac{3}{2}}$, with $0<\Gamma^{2D}\leq\Gamma^{*}$.
Figure 10: (a) The $(L_{0},\Gamma^{2D})$-parameter map of the classical 2D vertex model [1]. Cells are rigid in the white region, becoming floppy for $L_{0}>s_{6}\approx 3.72$. No solutions exist in the hatched region. (b) Area $A_{h}$ and perimeter $L_{h}$, and (c) pressure $P_{h}^{2D}$ and tension $T_{h}^{2D}$, are plotted as functions of $L_{0}$ for $\Gamma=$ 0.24; when $L_{0}=1$, $A_{h},L_{h}$ take the base case values in table 1.

### v2

@@ III.1 Regular hexagonal cells: parameter variation
$\Gamma_{A}$ controls the strength of cell-cell adhesion, so that increasing $\Gamma_{A}$ acts to increase lateral face area. Accordingly, decreasing $\Gamma_{A}$ penalises contact to the extent that $H$ falls to zero at $\Gamma_{A}\approx-0.38$ (Figure 4). While the cell height monotonically increases with $\Gamma_{A}$, there is a maximum in apical area and perimeter at $\Gamma_{A}\approx-0.25$. Strong adhesion elongates cells but also drives them into a floppy regime. This can be understood from the terms $\tfrac{1}{2}\Gamma_{A}H-2\Gamma_{L}L_{0}$ in $T^{2D}$ in (14c): increasing $\Gamma_{A}$ has an effect roughly equivalent to increasing $L_{0}$, which as already been seen to induce loss of rigidity. Once in the floppy regime, the asymptote (51) is recovered, with $H\propto-\Gamma_{A}$, $A\propto-\Gamma_{A}^{-1}$ and $L\propto-\Gamma_{A}^{0}$, consistent with very strong adhesion driving the formation of columnar cells. We will show below how columnar cells can also remain rigid for large $\Gamma_{A}$.
@@
Figure 4: Impact of varying $\Gamma_{A}$, presented following the format of Fig. 2. (b) The baseline cell shape at $\Gamma_{A}=-0.5$ is shown in grey (solid symbol); cells for $\Gamma_{A}={0.25,-1.0}$ are marked by open symbols. In all panels, $\Gamma_{a}=\Gamma_{L}=0.1,L_{0}=2.0$ and $a_{0}=1$. The dashed lines in (a,c) show the asymptote (51) as $\Gamma_{A}\rightarrow\infty$.
@@
Figure 5: Impact of varying $\Gamma_{L}$, presented following the format of Fig. 2. (b) The baseline cell shape at $\Gamma_{L}=0.1$ is shown in grey (solid symbol); cells for $\Gamma_{L}={0.025,0.25}$ are marked by open symbols. In all panels, $\Gamma_{a}=0.1,\Gamma_{A}=-0.5,L_{0}=2.0$ and $a_{0}=1$. The dashed lines show the asymptote as $\Gamma_{L}\rightarrow\infty$ as calculated in (48).
Figure 13 (Appendix I below) shows the impact of varying the target surface area $a_{0}$ from unity with other parameters held at their baseline values. Increasing $a_{0}$ takes the system into a floppy state, while causing cells to elongate.
@@
The transitions described above are summarized in Fig. 6, which shows six 2D slices through parameter space, for hexagonal cells. In comparison to the standard 2D model, there are multiple paths in parameter space along which the system loses rigidity, or along which rigid solutions become unphysical. The $(L_{0},\Gamma_{L})$-plane (Fig. 6c) is the most direct analogue of the standard 2D model (Fig. 10), but illustrates significant variation in the value of $L_{0}$ at which the rigidity transition takes place. Figure 6(a) offers an example where rigid solutions are lost either because the apical area vanishes (by decreasing $L_{0}$ from the baseline value, for fixed $\Gamma_{A}$), or because of the termination of a steady solution branch at a saddle-node bifurcation (by simultaneous reduction of $L_{0}$ and $\Gamma_{A}$), at which the eigenvalue corresponding to the dilational mode of the Hessian vanishes. Figure 6(e) shows how variation of both $\Gamma_{A}$ and $\Gamma_{a}$ can lead to loss of rigidity (increasing $\Gamma_{A}$ from the baseline case), flattening of the cell (reduction of $\Gamma_{A}$) or cell elongation (via simultaneous increase of $\Gamma_{A}$ and $\Gamma_{a}$). Fig. 6(d,e) demonstrates distinct floppy states, either squamous (low $\Gamma_{a}$) or columnar (high $\Gamma_{a}$); these are reflected in the two branches of the rigid/floppy boundary in Fig. 6(a,b). Figure 14 (Appendix I below) maps out variations of $a_{0}$ versus $\Gamma_{a}$, $\Gamma_{A}$, $\Gamma_{L}$ and $L_{0}$. For large $\Gamma_{a}$ or large $\Gamma_{L}$ (constraining the total area or perimeter), rigid solutions are confined to a window of $a_{0}$ values, demonstrating that rigidity can be lost by a reduction in $a_{0}$.
@@
We used asymptotic approaches (Appendix E) to identify the limiting location of the rigid/floppy transition, such as the limit $\Gamma_{a}\approx\Gamma_{L}L_{0}$ (52) for large $\Gamma_{A}$ (Fig. 6a,e,f), (45) for large $\Gamma_{a}$ (Fig. 6b,d,e and 14a) and (48) for large $\Gamma_{L}$ (Fig. 6c,d,f and 14c). The boundary at which rigid cells become strongly elongated in Fig. 6(b) is $L_{0}\approx-\sqrt{3}/(48\Gamma_{L})$ (46b)3. Similarly, rigid cells become strongly flattened at the threshold (49) for large $\Gamma_{L}$ (Fig. 6(f); the boundary turns out to be almost independent of $\Gamma_{L}$) and columnar at $\Gamma_{a}\approx-L_{0}^{3}/192$ (50) (Fig. 6c). These limiting approximations reflect dominant physical balances. Cells can remain rigid for large $\Gamma_{A}$ when $\Gamma_{a}$ is also large (Fig. 6e), with the associated tensions balancing to determine $H\approx\sqrt{3}\Gamma_{A}^{2}/(96\Gamma_{a}^{2})$ (53). In none of these limits does the 3D model recover results from the classical 2D model.
@@
Figure 6: Cross-sections of $(L_{0},\Gamma_{a},-\Gamma_{A},\Gamma_{L})$-parameter space, intersecting the baseline case (black circle). In each plot, only two parameters are varied. Hatching shows regions where there are no physical solutions. Grey regions show the floppy regime. Blue lines show where $T^{2D}=0$, corresponding to a rigidity transition; blue dashed lines show asymptotes at $\Gamma_{a}\rightarrow\infty,\Gamma_{A}\rightarrow$-$\infty$ and $\Gamma_{L}\rightarrow\infty$, calculated using equations (45), (52) and (48) respectively. Black dashed lines show asymptotes marking where $H\rightarrow 0$ (for large $\Gamma_{L}$, using (49), in (f)) or where $H\rightarrow\infty$ (for large $\Gamma_{a}$, using (46b), in (b, e); and for large $\Gamma_{L}$, using (50, in (c)).
The slices of parameter space in Figs 6 and 14 do not access a further limiting case, namely that of constrained cell volume, for which $\Gamma_{a}$, $|\Gamma_{A}|$ and $\Gamma_{L}$ are all small in magnitude. We show in Appendix E how a sharp squamous to columnar transition can arise as the ratio $\Gamma_{A}/\Gamma_{a}$ varies when $a_{0}>2^{2/3}3^{7/6}\approx 5.71$ when $\Gamma_{L}=0$, mirroring a finding by [31] using a closely related model. Fig. 11 below illustrates how, for $\Gamma_{L}>0$, cells can become floppy through a rapid but continuous transition.
@@ III.2 Regular hexagonal cells: external loading
The slopes of the volume curves in Fig. 7(c) illustrate the cell bulk modulus under in-plane and isotropic compression. The full stiffness matrix is evaluated in Appendix H. Second derivatives of the energy in (63) illustrate how the coupling between height, perimeter and area variations in the present 3D formulation contribute to stiffness. The first derivatives of the energy in (63) illustrate the contribution of prestress.
@@ III.3 Disordered Monolayers
Figure 8: Distributions of cell height perturbations $H^{(1)}$ evaluted using (44), for three different monolayers using baseline parameters, $L_{0}=2,\Gamma_{a}=0.1,\Gamma_{A}=-0.5,\Gamma_{L}=0.1$ and $a_{0}=1$.
@@
Figure 9: Simulations of a monolayer with $N_{c}=500$ cells and $N_{v}=1080$ apical vertices, successively relaxed through values of $L_{0}$. (a) Distribution of the shape parameter ${L_{h}}/{\sqrt{A_{h}}}$ for varying $L_{0}$ with $\Gamma_{a}=0.1,\Gamma_{A}=-0.5,\Gamma_{L}=0.1$ and $a_{0}=1$, coloured by the number of apical edges of each cell. White circles indicate outliers and the whiskers show the data extent excluding outliers, while the box shows the interquartile range. The red dots indicate the value of the shape parameter for an array of hexagons, which here is above the value at transition of $s_{6}\approx 3.72$. Thresholds $s_{4}=4$ and $s_{5}\approx 3.81$ (see (18)) are shown with dashed lines. (b) The eigenvalues of the Hessian for the monolayers show in (a), for $L_{0}=\{2.85,2.9,2.95,3.0\}$. (c) Number of connected components $N_{cc}^{\epsilon}$ in each monolayer (blue squares, left axis) and number of vertices in the largest connected component $N^{\varepsilon}_{\max}$ (red diamonds, right axis) as $L_{0}$ is varied. (d) Monolayers for $L_{0}=\{2.88,2.89,2.9,2.95\}$ showing edges with tension above a threshold value of $\varepsilon=0.0005$ in red.
@@ IV Discussion
We have shown how a planar epithelium, modelled using a mechanical energy $U$ that reflects the 3D structure of cells, can be represented in a 2D formulation. To achieve this, we introduced a constraint (42) that restricted height variations between neighbouring cells. A global force balance was used to determine the height $H$ of the cell population. This, together with the involvement of an energy that penalises deviations of the total cell area from a target value, leads to a 2D model that is distinct from the classical formulation in which apical area $A$ and perimeter $L$ contribute separately to $U$. The coupling of $A$, $H$ and $L$ leads to a more intricately-structured parameter space (Figs 6, 14) that reveals multiple pathways to rigidity transitions, regulated by variation of five independent dimensionless constitutive parameters. This suggests that a monolayer changing state, for example in a wound-healing response, need not rely wholly on the cortical actin structures at the cell edge. This is consistent with prior models [11] demonstrating regulation of rigidity via variations of the bulk target area through, for example, the parameter $a_{0}^{*}/(V_{0}^{*})^{2/3}$. Like [31], our model predicts a sharp squamous to columnar transition arising when the cell volume is constrained, provided $a_{0}^{*}/(V_{0}^{*})^{2/3}$ is sufficiently large. Our findings rest on an assumption of apical-basal symmetry, so are likely to be of less relevance to epithelia that adhere strongly to a basement membrane, which can be expected to promote monolayer rigidity [30].
@@ Appendix A Incidence and adjacency matrices
Assigning orientations to all quantities, signed incidence matrices $\mathsf{A}$, $\mathsf{B}$ and $\mathsf{C}$ with components $C_{hi}=\{\mathsf{C}\}_{hi}$, $B_{ij}=\{\mathsf{B}\}_{ij}$ and $A_{jk}=\{\mathsf{A}\}_{jk}$ define topological arrangements. Specifically, an edge orientation is defined by a tangent with magnitude equal to the edge length; a face orientation is defined by a normal with magnitude equal to the face area; $C_{hi}=1$ (1-1) if the normal to face $i$ points into (out of) cell $h$; $B_{ij}=1$ (1-1) if edge $j$ bounds face $i$ and is congruent (or anticongruent) to its orientation in a right-handed sense; $A_{jk}=1$ (1-1) if edge $j$ points into (out of) vertex $k$. The components of $\mathsf{A}$, $\mathsf{B}$ and $\mathsf{C}$ are zero otherwise.
@@
The factor $\tfrac{1}{6}$ in $\mathsf{D}^{\mathrm{cv}}$ accounts for the fact that each vertex $k$ of cell $h$ neighbours three edges of the cell and each edge bounds two faces of the cell. For the prismatic cells illustrated Fig. 1, distinguishing apical (++), lateral $(l)$ and basal ()(-) objects, the adjacency matrices inherit the structure of the incidence matrices via
@@ Appendix B Geometric identities
[…]
@@
Rigid hexagons satisfy (16a,b) (with $a_{0}=1$); floppy solutions satisfy $A_{h}=1$, $L_{h}=L_{0}^{2D}$, as illustrated in Fig. 10(b,c) (replicating [1], for example), which can be compared to figures 2. The parameter space map in Fig. 10(a) can be compared to Fig. 6(c). The rigid-floppy transition lies at $L_{0}^{2D}=s_{6}\approx 3.72$. Rigid solutions collapse to zero area along $L_{0}^{2D}=0$ for $\Gamma^{2D}>\Gamma^{*}\equiv\sqrt{3}/12$. Rigid solution branches terminate via a saddle-node bifurcation at $L_{0}^{2D}=-({8\sqrt{2}}/(3\Gamma))\left(\Gamma^{*}-\Gamma^{2D}\right)^{\tfrac{3}{2}}$, with $0<\Gamma^{2D}\leq\Gamma^{*}$.
Figure 10: (a) The $(L_{0},\Gamma^{2D})$-parameter map of the classical 2D vertex model [1]. Cells are rigid in the white region, becoming floppy for $L_{0}>s_{6}\approx 3.72$. No solutions exist in the hatched region. (b) Area $A_{h}$ and perimeter $L_{h}$, and (c) pressure $P_{h}^{2D}$ and tension $T_{h}^{2D}$, are plotted as functions of $L_{0}$ for $\Gamma=$ 0.24; when $L_{0}=1$, $A_{h},L_{h}$ take the baseline values in table 1.